OpenStudy (anonymous):

If the sin 75° is 0.966, then the cos _____° is 0.966. @Compassionate

OpenStudy (anonymous):

if Cos(x) = sin(y) what connection does x have with y ?!

OpenStudy (anonymous):

Just says that numerical answers are expected

OpenStudy (compassionate):

No, Newman, x and y are ________ to one another?

OpenStudy (anonymous):

congruent ?

OpenStudy (compassionate):

They're complementary.

OpenStudy (anonymous):

yea that to

OpenStudy (compassionate):

If the sin 75° is 0.966, then the cos _____° is 0.966. So we know sin is 75, and sin and cos will ad up to 90 degrees. Therefore?

OpenStudy (anonymous):

75

OpenStudy (compassionate):

If sine and cosine will add up to 90 degrees because they are complementary, so 75 + x = 90, what is x?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

if 10 + x = 15 how much is x ? :P

OpenStudy (compassionate):

Come on! We can do it!

OpenStudy (anonymous):

♪ ♫ we can do it.. we we can do it ♪ ♫

OpenStudy (anonymous):

sephy help

OpenStudy (compassionate):

We are helping you. This is basic. We're not even asking you geometry, it's basic arithmetic.

OpenStudy (anonymous):

all you need to do is \[\theta=\cos^{-1} 0.966\] =14.98

OpenStudy (anonymous):

Idk nig nog

OpenStudy (compassionate):

90 - 75

OpenStudy (anonymous):

Lol sorry I wasn't paying attention.. I was playing gta... 15

OpenStudy (compassionate):

If you're not paying attention, then why use OpenStudy?

OpenStudy (anonymous):

I was busy

OpenStudy (shaw74):

thank you soooo much I really understand this now! it was so easy!

OpenStudy (anonymous):

wait so its 15, right?

OpenStudy (anonymous):

WOW BUSY PLAYING GTA